Bbxblue23:

how do i do thiss "write the equation of a parabola in vertex form where the vertex is a maximum located in quadrant 3 "

balling4days:

does it have a graph or something

Bbxblue23:

no its just the question like

@balling4days wrote:
does it have a graph or something
that

R3B3LMARINE:

y=|x|-3

Bbxblue23:

@r3b3lmarine wrote:
y=|x|-3
is that the answer or the formula??

R3B3LMARINE:

y=|x|-5 is an equation. It should get you the correct answer.

Bbxblue23:

idk how to do this tho, like how to solve it

@r3b3lmarine wrote:
y=|x|-5 is an equation. It should get you the correct answer.

R3B3LMARINE:

@bbxblue23 wrote:
idk how to do this tho, like how to solve it
@r3b3lmarine wrote:
y=|x|-5 is an equation. It should get you the correct answer.
You asked for the equation. That is the equation.

Bbxblue23:

@r3b3lmarine wrote:
@bbxblue23 wrote:
idk how to do this tho, like how to solve it
@r3b3lmarine wrote:
y=|x|-5 is an equation. It should get you the correct answer.
You asked for the equation. That is the equation.
ight ty

R3B3LMARINE:

No problem. I am sorry if you do not get it, I had a hard time with Parabolas, and they are annoying.

Bbxblue23:

@r3b3lmarine wrote:
No problem. I am sorry if you do not get it, I had a hard time with Parabolas, and they are annoying.
another question, doyou know how to solve y=x^2-14+40 using the x= -b over 2a??

R3B3LMARINE:

@bbxblue23 wrote:
@r3b3lmarine wrote:
No problem. I am sorry if you do not get it, I had a hard time with Parabolas, and they are annoying.
another question, doyou know how to solve y=x^2-14+40 using the x= -b over 2a??
That I do not know. I am sorry.